Suppose an airline accepted 12 reservations for a commuter plane with 10 seats. They
know that 7 reservations went to regular commuters who will show up for sure. The
other 5 passengers will show up with a 50% chance, independently of each other.
(a) Find the probability that the flight will be overbooked.
(b) Find the probability that there will be empty seats.
1
Expert's answer
2020-05-19T19:06:14-0400

(a) Find the probability that the flight will be overbooked.

"P(X=4)+P(X=5)=" C5,4(0.5)4(1−0.5)5−4+C5,5(0.5)5(1−0.5)5−5=(5+1)(0.5)5=0.1875

(b) Find the probability that there will be empty seats.

"P(X=0)+P(X=1)+P(X=2)=" C5,0(0.5)0(1−0.5)5−0+C5,1(0.5)1(1−0.5)5−1+C5,2(0.5)2(1−0.5)5−2=(1+5+10)(0.5)5=0.5